### v1

Abstract
Decentralized social media platforms like Bluesky Social (Bluesky) have made it possible to publicly disclose some user behaviors with millisecond-level precision. Embracing Bluesky’s principles of open-source and open-data, we present the first collection of the temporal dynamics of user-driven social interactions. BlueTempNet integrates multiple types of networks into a single multi-network, including user-to-user interactions (following and blocking users) and user-to-community interactions (creating and joining communities). Communities are user-formed groups in custom Feeds, where users subscribe to posts aligned with their interests. Following Bluesky’s public data policy, we collect existing Bluesky Feeds, including the users who liked and generated these Feeds, and provide tools to gather users’ social interactions within a date range. This data-collection strategy captures past user behaviors and supports the future data collection of user behavior.
[…]
DATA TYPE/LOCATION Network Graph; Bluesky Social (Social Media Platform)
{IEEEkeywords}
Blocking and Group Sanction, Bluesky Social, Community Analysis, Decentralized Social Media, Multi-network Analysis, Social Networks, Temporal Graph Mining, User Behavior Analysis
1 Background
Understanding the evolution of user behavior through complex interactions in online social networks has been an important topic in computational social science [21, 17, 7, 9, 3, 11], and computer science [28, 16, 22, 26]. However, researchers face two challenges in studying user behavior on social media platforms: (1) platforms typically design recommendations to target users, making it difficult to isolate their effects on user behavior; and (2) platforms often limit access to detailed timing and comprehensive records of how users interact. Table 1 categorizes and compares how eight social media platforms manage key aspects of users’ social interactions, including following (establishing one-way connections and receiving updates from connected users), blocking (preventing visibility and engagements between users), creating (generating a new community), and joining (subscribing and receiving a community’s suggested content).
Table 1: Comparing accessible social interactions of users on social media platforms via APIs. The symbols used are: $\times$ for complete exclusion, $\circ$ for imperfect coverage (i.e., no timestamp), and $\bullet$ for perfect coverage.
Platforms User-to-User User-to-Community
Follow Block Create Join
Twitter [27, 19] $\circ$ $\times$ $\times$ $\times$
Facebook [23] $\circ$ $\times$ $\times$ $\times$
Instagram [32] $\circ$ $\times$ $\times$ $\times$
LinkedIn [6, 30] $\times$ $\times$ $\times$ $\times$
Reddit [2, 10] $\times$ $\times$ $\circ$ $\circ$
TikTok [5] $\times$ $\times$ $\times$ $\times$
Mastodon [20, 34] $\circ$ $\times$ $\bullet$ $\bullet$
Bluesky Social $\bullet$ $\bullet$ $\bullet$ $\bullet$
Interest in alternatives to the central control of traditional social media platforms has led to a substantial migration of users to decentralized platforms [14]. Here, we focus on Bluesky [^1] https://bsky.social/about, a social media platform currently boasting over 5 million active users, where data are distributed and managed across multiple providers. Despite its decentralized architecture, Bluesky offers a seamless user experience comparable to that of centralized platforms. This is achieved through interoperability facilitated by its protocol, ATproto [18]. Bluesky features a default reverse chronological timeline and allows users to add customized recommendation algorithms created by other users. This user-driven design facilitates the study of user behavior without the potential control exerted by platforms [25].
Following Bluesky’s public data policy [^2] https://blueskyweb.zendesk.com/hc/en-us/articles/15835264007693-Data-Privacy and leveraging the advanced data accessibility and accountability offered through the Personal Data Server (PDS) [^3] https://en.wikipedia.org/wiki/Personal_data_service model, we collected the temporal dynamics of social interactions among over 150,000 public users. Our dataset, BlueTempNet, includes two categories of users’ social interactions with millisecond-level timestamps. Figure 1 illustrates how these interactions are integrated into a multi-network framework, bridging various types of nodes and edges while allowing multiple directional connections from the same nodes. BlueTempNet offers a unique opportunity for comprehensive analysis, spanning from individual actor networks to the platform ecosystem over time.
Figure 1: Overview of BlueTempNet, a multi-network integrating two categories of users’ social interactions such as user-to-user (following, blocking) and user-to-community (creating, joining), all timestamped at $t_{i}$.
The key contributions of our work are:
  • A novel dataset, BlueTempNet, captures the temporal dynamics of users’ social interactions, independent of any platform-driven recommendation algorithms.
  • BlueTempNet integrates key behavioral aspects of users’ social interactions into a multi-graph with both user-to-user and user-to-community networks.
  • BlueTempNet contains longitudinal data, which is essential for studying the development and sustainability of Bluesky since the introduction of custom Feeds.
Figure 2: An example of a user’s interactions on Bluesky. The left panel shows user-to-community interactions, including joining a community (purple dashed line) and a community creation (green dashed line). The right panel illustrates user-to-user interactions, the change in profiles after following (blue arrow) and blocking (red arrow) another user.
2 PLATFORM INTERFACE
To gain a clearer understanding of Bluesky’s user experience, Figure 2 illustrates social interactions using the corresponding authors’ account as an example. On the right side of the figure, the follow action represents a user’s interest in another user, allowing them to receive updates and posts in their default timeline. The block feature enhances privacy and safety by ensuring that users who have blocked each other cannot view or engage with each other’s posts, making them invisible to both sides. On the left side of the figure, there is a list of Feeds, which are dynamic content streams with customizable algorithms. Users join Feeds of interest, like philsky (Philosophy content) and Discover (Trending content), by clicking the heart icon below their display names and descriptions, an action known as Feed like. Each Feed is created by a user using ATproto’s Feed generator.
3 COLLECTION METHODS AND DESIGN
We used the Bluesky API provided by ATproto [^4] https://atproto.com/ protocol. As shown in Figure 3, there are three sequential steps in the data collection process: (1) collecting existing Feeds from Bluesky at the community level; (2) gathering information about users who generated or liked these Feeds at the user-to-community level; and (3) compiling relationships between users by examining if they blocked or followed each other at the user-to-user level.
4 RECORDS AND STORAGE
4.1 Storage Principles
In accordance with FAIR (Findability, Accessibility, Interoperability, and Reuse of digital assets) principles [8], we took measures to ensure that the data we utilized or curated are devoid of personally identifiable information. Any text-level information about users’ profiles or their posts is not covered in the dataset for privacy measures. Throughout our analysis, we anonymized user identifiers with a MongoDB database through field-level encryption.
4.2 Scope of Records
The versatile data collection framework accommodates different periods of activity established for the research. This study focuses on the year following the introduction of custom Feeds on Bluesky [^5] https://bsky.social/about/blog/7-27-2023-custom-feeds, from May 11, 2023, to May 11, 2024. Based on the time of our data collection.
Figure 3: An illustration of our data collection pipeline on Bluesky is presented in three levels. Each stage of collection captures the timestamp of user behavior with millisecond-level precision in UTC format.
4.3 Definition of User
Though Bluesky does not formally categorize users, we defined two types based on their relationships with Feeds. This simplifies understanding and enhances the hierarchical structure of our proposed multi-graph. These two user types are defined as follows:
  • Feed Creators: Users who create Feeds using the Feed generator. These users can provide content for others by proposing their own recommendation algorithms.
  • Feed Members: Users who join Feeds by liking them. These users primarily consume the curated content provided by the Feed’s recommendation algorithm.
Table 2 presents statistics of Bluesky, including the number of users (Feed creators and members), the total Feeds, and the engagement counts from Feed members. In order to integrate their connections into a multi-graph through their shared Feeds, we only focused on the interactions among the Feed creators and Feed members
Table 2: Summary of the user and Feed samples identified in Bluesky. Active users and Feeds are set at the time of data collection (May 11, 2024).
Types Number
Active Users 5,659,340
Active Feeds 39,968
Feed Creators (unique active users who have generated at least one active Feed) 17,617
Feed Members (unique active users who have liked at least one active Feed) 147,577
Total Like Count on Feeds by Feed Members 322,930
Total Block Count of Feed Members 2,677,622
Total Follow Count of Feed Members 24,482,447
4.4 Dimensions of Records
Figure 3 shows our data collection process on Bluesky. Based on users, Feeds, and their follow and block relationships at each timestamp of social interaction, we defines three dimensions of user and community interactions.
  1. 1.
    Feed Creator Interaction Network: A signed graph representing follows (positive edges) and blocks (negative edges) among Feed creators, denoted as $G_{\mathcal{C}}=\{\mathcal{C},\mathcal{E}^{+},\mathcal{E}^{-}\}$, where $\mathcal{C}$ is the set of Feed creators.
  2. 2.
    Feed Member Interaction Network: A signed graph representing follows (positive edges) and blocks (negative edges) among users, denoted as $G_{\mathcal{M}}=\{\mathcal{M},\mathcal{E}^{+},\mathcal{E}^{-}\}$, where $\mathcal{M}$ is the set of Feed members.
  3. 3.
    Community Interaction Network: An affiliation graph connecting communities of users with Feeds, denoted as $G_{\mathcal{A}}=\{\mathcal{V},\mathcal{E}\}$, where $\mathcal{V}=\{\mathcal{C},\mathcal{M},\mathcal{F}\}$ represents three types of nodes (creators, members, and Feeds), and =$\mathcal{E}=\{\mathcal{G},\mathcal{L}\}$ represents two types of undirected edges (Feed generation and Feed likes).
Following and blocking are not mutually exclusive actions. A user can follow others initially and then block them later. For these cases, we prioritize the blocking interaction by retaining the negative edges (blocking) and removing the positive edges (following). Based on this, we present the statistics for the three dimensions of the dataset: (1) Table 3 shows the Feed creator interaction network. The graph contains 349 isolated nodes and 485 node pairs with both negative and positive edges. (2) Table 4 presents the Feed member interaction network. The graph includes 34 isolated nodes and 5,502 node pairs with both negative and positive edges. (3) Table 5 describes the community interaction network. In the affiliation graph, we identified 8,820 users who act as both Feed creators and Feed members. These users are represented as two nodes: one as a member node and the other as a creator node.
Table 3: First dimension – statistics of the signed graph $G_{\mathcal{C}}$, showing blocking and following among the Feed creators identified in Table 2.
Types Number
Node (Feed Creators) 17,617
Negative Edge (Blocking among Feed Creators) 26,992
Positive Edge (Following among Feed Creators) 292,331
Table 4: Second dimension – statistics of the signed graph $G_{\mathcal{M}}$, showing blocking and following among the Feed members identified in Table 2.
Types Number
Node (Feed Members) 147,577
Negative Edge (Blocking among Feed Members) 532,614
Positive Edge (Following among Feed Members) 5,766,570
Table 5: Third dimension – statistics of the affiliation graph $G_{\mathcal{A}}$, showing the relations across Feeds, creators, and members in Tables 3 and 4.
Types Number
Member Node (Feed Members) 147,577
Creator Node (Feed Creators) 17,617
Feed Node (Active Feeds) 39,968
Join Edge (Liking Feeds) 322,930
Create Edge (Genearting Feeds) 39,968
5 INSIGHTS AND NOTES
5.1 Feed Distribution
Figure 4 shows the popular Feeds and the frequent terms used in the display name of Feeds. As the 64-character limit for Feed names will encourage users to choose words carefully, we assume users selectively choose words in display names to represent the interest of the Feed. By removing redundant words such as feed, tab, post, and test, our analysis revealed that a large number of Feeds are found to focus on topics such as art, media, music, japanese, ai, and nsfw. This indicates that Bluesky users have a focused preference for niche content.
Figure 4: Bar charts showing (A) top 20 most liked Feed’s display names and (B) top 20 most frequently used terms in Feed’s display names.
The popularity of Feed does not necessarily correlate with its prevalence in Bluesky. To examine the relationship between display names and the popularity of Feeds, we calculated term frequency vectors and computed Pearson correlation coefficients for each word. Figure 5 shows words with the highest and lowest correlations of words. Feeds with words like likes, discover, and hot were strongly associated with popularity. Community and social topics such as blacksky (Black community), ukraine (Ukraine-Russian war), academic (science, astronomy, etc), transqueer (LGBTQ), and adult content also showed positive correlations. likes was the highest term, indicating that liking is perceived as an important behavior among users for discovering information on Bluesky. Terms like art, gallery, and media were among the least popular despite their prevalence in Feeds’ display names. However, the correlation for these terms is very low (below $-0.005$), implying minimal impact on Feed’s popularity.
Summary (Communities) There is a notable contrast between the Feeds users prefer to create and those that achieve widespread popularity. On Bluesky, niche Feeds related to art and media are prevalent. However, Feeds that facilitate content discovery and highlight trending topics tend to garner the most number of likes from users.
Figure 5: Bar charts showing Pearson correlations between term frequency in Feed’s display names and the number of likes: (A) terms with top 20 correlations, (B) terms with bottom 20 correlations.
5.2 Temporal Dynamics
Figure 6 shows the temporal trends of three dimensions of social interactions in our dataset. We identified a sudden surge in both follows and blocks after Bluesky transitioned to an invitation-free platform on February 6, 2024. Among Feed creators, follows increased by 19.2% and blocks by 27.3%. Similarly, among Feed members, follows increased by 15.8% and blocks by 25.9%. After the transition, 44% of the Feeds were generated, and users liked 34.12% of them. We speculate that the larger increase in blocks compared to follows indicates a heightened need for users to filter out potentially unwanted interactions from new users and Feeds, especially after the platform became completely open.
Figure 7 illustrates the evolution of the degree distribution across three dimensions of our dataset. The average and standard deviation of in-degrees and out-degrees gradually increase over time, showing stabilization as time progresses. Despite the surge in users and interactions during the transition to an invitation-free platform, the average and standard deviation of interactions remain stable across all three dimensions. This indicates that Bluesky managed the influx of users and their interactions, maintaining stability and aligning with existing trends during its establishment.
Summary (Temporal Dynamics) After Bluesky’s official public launch, social interactions surged across all dimensions. The rate of user blocks increased more sharply compared to the rate of follows, and the creation of new Feeds outpaced the rate at which users joined them. Despite this surge in users and their interaction levels, Bluesky managed to stabilize the interactions to some extent before the transition to an invitation-free platform.
Figure 6: Temporal trends in the number of nodes (orange) and edges (blue) for three networks: $G_{\mathcal{C}}$ (Feed Creator Interaction), $G_{\mathcal{M}}$ (Feed Member Interaction), and $G_{\mathcal{A}}$ (Community Interaction). The red dotted line marks when Bluesky transitioned to an invitation-free platform.
Figure 7: Temporal trends in degree distributions for three networks: $G_{\mathcal{C}}$ (Feed Creator Interaction), $G_{\mathcal{M}}$ (Feed Member Interaction), and $G_{\mathcal{A}}$ (Community Interaction). Lines show average in-degree (blue) and out-degree (orange); shaded areas indicate standard deviation. The red dotted line marks when Bluesky transitioned to an invitation-free platform.
5.3 Network Analysis
We base our network analysis on the static graph from the last data collection on May 11, 2024. For simplicity, we focus on the largest component of each graph, encompassing 88% of nodes in $G_{\mathcal{C}}$, 94% of nodes in $G_{\mathcal{M}}$, and 86% of nodes in $G_{\mathcal{A}}$. To comprehensively view each network’s topology, we carefully selected applicable measures for the signed graph [4] and measures for the affiliation graph [24].
5.3.1 User-to-User (singed graph) Networks Analysis
Table 6 summarizes metrics for the signed graphs $G_{\mathcal{C}}$ and $G_{\mathcal{M}}$, highlighting five key aspects: (1) Positive assortativity is slightly negative in both graphs, indicating high-degree users tend to follow lower-degree users. (2) Negative degree assortativity is higher for $G_{\mathcal{C}}$ compared to $G_{\mathcal{M}}$, showing distinct blocking patterns between creators and members. (3) Positive reciprocity is higher than negative reciprocity in both graphs, suggesting stronger mutual following than blocking. (4) Both graphs exhibit similar positive clustering coefficients but low negative clustering coefficients, indicating that followers’ neighbors are more likely to be connected than blockers’ neighbors. (5) Structural balancedness is slightly higher in $G_{\mathcal{C}}$ than in $G_{\mathcal{M}}$, suggesting a greater prevalence of social balance, where an enemy’s enemy is considered a friend, among Feed creators.
Table 6: Global-level metrics of User-to-User networks are provided with values rounded to two decimal places. In each signed graph, path lengths are determined solely using positive edges, as blocking edges prevent visibility and interaction between users, thereby there is no flow of information.
Metric Description $G_{\mathcal{C}}$ $G_{\mathcal{M}}$
Diameter The greatest distance between any pair of nodes 12 11
Average Path Length The average shortest path between all pairs of nodes 3.18 3.19
Degree Density Proportion of actual edges compared to possible edges 1.07 $e^{-3}$ 2.97 $e^{-4}$
Positive Degree Assortativity Bias for nodes to follow neighbors with similar following degrees -0.13 -0.03
Negative Degree Assortativity Bias for nodes to block neighbors with similar blocking degrees 0.02 -0.15
Positive Reciprocity Likelihood of nodes to be mutually linked by following edges 0.46 0.50
Negative Reciprocity Likelihood of nodes to be mutually linked by blocking edges 0.07 0.06
Positive Clustering Coefficient Average of two adjacent nodes of a follower also follow each other 0.24 0.25
Negative Clustering Coefficient Average of two adjacent nodes of a blocker also block each other 0.03 0.03
Structural Balancedness Ratio of socially balanced triangles based on undirected edges 0.91 0.89
Table 7: Global-level metrics for User-to-Community networks where edges are undirected. Values for each metric are rounded to two decimal places.
Metric Description $G_{\mathcal{A}}$ $P_{\mathcal{A}}$
Diameter The greatest distance between any pair of nodes 24 11
Average Path Length The average shortest path between all pairs of nodes 60.49 3.06
Degree Density Proportion of actual edges compared to possible edges 1.87 $e^{-5}$ 1.07 $e^{-2}$
Degree Assortativity Nodes to connect with similar degrees of neighborhood -0.16 0.50
Clustering Coefficient Average of two neighbors of a node are also neighbors to each other 0.80 0.92
5.3.2 User-to-Community (affiliation graph) Networks Analysis
To analyze higher-order relationships between communities, we define a line graph $P_{\mathcal{A}}=\{\mathcal{F},\mathcal{E},\mathcal{W}\}$ by projecting the affiliation graph $G_{\mathcal{A}}$ as follows:
$\mathcal{E}$ =$=\left\{(f_{i},f_{j})\mid U(f_{i})\cap U(f_{j})\neq\emptyset\right\}$ (1)
$w(f_{i},f_{j})$ $=\left|U(f_{i})\cap U(f_{j})\right|,\>\text{with}\>w\in\mathcal{W}$
where $U(f_{i})$ denotes the set of any users adjacent to one of Feeds $f_{i}\in\mathcal{F}$ from $G_{\mathcal{A}}$. Two nodes $f_{i}$ and $f_{j}$ are linked if they share at least one user. The weight of an edge $w(f_{i},f_{j})$ is the number of shared users between the Feeds.
Table 7 compares the affiliation graph $G_{\mathcal{A}}$ with its projected line graph $P_{\mathcal{A}}$, revealing three key differences: (1) the degree density is higher in $P_{\mathcal{A}}$ than in $G_{\mathcal{A}}$. (2) $P_{\mathcal{A}}$ exhibits a short average path length and a high clustering coefficient, which is close to the maximum, indicating the properties of a small-world network [31]. (3) the degree assortativity for $G_{\mathcal{A}}$ is negative, whereas it is positive in $P_{\mathcal{A}}$, showing that Feeds with more members tend to connect together.
Figure 8 illustrates the community analysis of the projected line graph $P_{\mathcal{A}}$, which are clustered using the Leiden Algorithm [29], with annotations derived from Feeds’ descriptions. The clusters reveal a variety of groups representing different languages and regions, such as German, Japanese, Spanish, and English groups on Bluesky. Additionally, the network features clusters centered on political topics, like the Ukraine-Russia war, alongside more casual feeds, with a notable prevalence of art-related communities. As discussed during network analysis, the relationships between Feeds exhibit small-world properties. Further exploration could uncover tightly knit community groups, including higher-order analyses to identify strongly connected clusters and examine their characteristics with respect to language, political geography, and relevant social events.
Summary (Network Analysis) Followings for both creators and members exhibit negative degree assortativity, connecting to nodes with lower degrees. Members also tend to make blockers more connected, with higher structural balancedness highlighting distinct user roles. In community analysis, the projection of the affiliation graph reveals small-world properties, particularly with prominent art-related communities. Studying higher-order relations using the projected line graph appears promising for understanding how the platform maintains its ecosystem between communities.
Figure 8: A line graph of Feeds $P_{\mathcal{A}}$ is visualized using ForceAtlas2. A total of 770 clusters are identified by the Leiden Algorithm. The top 8 clusters are color-coded; others are gray. Each cluster’s interests are shown by word clouds, which aggregate the Feeds’ self-descriptions in each cluster.
5.4 LIMITATIONS
As Bluesky is still in development, user engagement may not be as consistent as on established platforms like Twitter/X. Consequently, usage patterns might be unstable and subject to changes due to platform updates, such as the end of a public open beta, or external factors. Additionally, users on Bluesky tend to focus on topics like art and AI, which may not generalize our findings to other social media platforms. Among many social interactions on Bluesky, we focused on the networking behavior among users, excluding activities related to posts, such as replying, liking, and reposting.
5.5 FUTURE WORK
We will examine the evolutionary aspects of rumor and misinformation spread [13, 12], the role of blocking mechanisms [9, 11], and platform-specific characteristics that promote sustainability of social media platforms [15, 14]. Testing the causal effects of user attributes and behaviors, along with evaluating the predictive power of social networks, is crucial for modeling future user behavior [33, 22]. Lastly, comparing social networks across diverse online environments like Epinions, Slashdot, and Wikipedia could reveal persistent interaction patterns in user behaviors that extend beyond typical social media platforms [1, 7, 22]
6 Source Code and Scripts
The scripts and dataset used for crawling Bluesky will be publicly available in our DOI [^6] https://ieee-dataport.org/documents/bluetempnet-temporal-multi-network-dataset-social-interactions-bluesky-social. The proposed graphs are saved in GEXF [^7] https://gephi.org/users/supported-graph-formats/ format and named as follows:
  • graph_dimension1.gexf: This file is the Feed member interaction network, where an edge has attributes sign and time and a node represents Feed member.
  • graph_dimension2.gexf: This file is the Feed creator interaction network, where an edge has attributes sign and time, and a node represents Feed creator.
  • graph_dimension3.gexf: This file contains the community interaction network. Each node has a type attribute, which can be one of the following: member, creator, or Feed. Each edge includes a time attribute.
  • multi_graph.gexf: This file is the multi-network integrating the three-dimensional graphs. For the sake of simplicity, undirected edges from the community interaction network are converted to bidirectional edges.
Each timestamp in the edge data is formatted in ISO 8601 (e.g., 2023-09-22T09:32:17.974Z) based on UTC.
Data collection scripts are written in Python 3 using NetworkX version 2.6.3. These scripts are named according to each dataset (e.g., graph_dimension1.py). While user IDs remain anonymous, we provide some public user activities and information on the collected Feeds as follows:
  • user_metadata.csv
    1. 1.
      Anonymized ID (decoded after ID request review)
    2. 2.
      Mapping between Anonymized ID and node ID
    3. 3.
      Number of Followers
    4. 4.
      Number of Following
    5. 5.
      Number of Posts
  • feed_metadata.csv
    1. 1.
      URI (a string of characters used to uniquely identify a Feed on the internet)
    2. 2.
      Mapping between URI and node ID
    3. 3.
      Display Name of Feed
    4. 4.
      Description of Feed
    5. 5.
      Creator of Feed
    6. 6.
      Number of Likes on Feed
ACKNOWLEDGMENTS AND INTERESTS
This work was funded by Office of Naval Research, under Award No. N00014-21-1-4002. Opinions, interpretations, conclusions, and recommendations within this article are solely those of the authors.